You invested 3000 in the stock market one year ago today to do the investment is value 3570. What return did you earn? What return would you suffer next year for the investment to be valued at the original 3000?
Answer #1

Invested Amount = P = $3000

Value after 1 year = F = $3570

Let the Return be r1

=> F = P(1+r1)

=> r1 = F/P - 1 = 3570/3000 - 1 = 0.19 = 19%

Let the return next year be r2

Value at end of Year 2 = X = $3000

=> X = F(1+r2)

=> r2 = X/F - 1 = 3000/3570 - 1 = -0.1597 or -15.96%